I run a small café. I hire one helper at Rs.180,000 per year, pay annual rent of Rs.150,000 for the hotel and spend Rs. 100,000 per year on materials. I had to put in Rs.40,000 of my own funds in equipment (five tables, ten chairs and so forth) that could have earned me Rs.4,000 per year if alternatively invested. I have been offered Rs.250,000 per year to work for a competitor. I estimate my entrepreneurial talents are worth Rs.30,000 per year. Total annual revenue from restaurant sales is Rs.720,000. Calculate the accounting profit and economic profit for my cafe.

 


1
Expert's answer
2021-03-30T07:17:21-0400

Accounting profit= total revenue - explicit costs

explicit costs = 180 000 + 150 000 + 100 000 = 430 000

=720000-430000

=Rs290, 000


Economic profit= total revenue - (explicit costs + implicit costs)

explicit cost = 430 000

implicit cost = 40, 000

=720000-(430000 +40000)

=Rs250, 000